# A single-layer framework of variational tensor network states

### v1

@@ III.1 The Heisenberg model on a square lattice
Figure 4: The PEPS ansatz used in this work to study the square lattice Heisenberg model. Dashed lines separate the $2\times 2$ unit cells.
The estimated ground-state energy is plotted in Fig. 5 as a function of $D$. It shows that the obtained energy systematically decreases as $D$ increases, verifying the variational nature of the framework. When $D=9$, we obtain an estimation about $E_{g}=-0.6693896$, which is already very close to the value $E_{g}=-0.6694421$ in quantum Monte Carlo simulation HM-MC . If we make a further power-law fitting of $E_{g}(1/D)$, then in the large-$D$ limit we obtain an extrapolation about $E_{g}=-0.66941$, which deviates from the quantum Monte Carlo estimation only by the order of $10^{-5}$.
Figure 5: The obtained ground-state energy $E_{g}$ of the square lattice Heisenberg model. (a) $E_{g}$ as a function of bond dimension $D$, with $D$ ranging from 2 to 9. (b) A power-law fitting of data $E_{g}(1/D)$. The extrapolation to the large-$D$ limit gives $E_{g}=-0.66941$. A quantum Monte Carlo estimation from Ref. HM-MC , $E_{g}=-0.6694421$, is also shown as red solid.
As mentioned before, the ground state of this model is antiferromagnetically long-range ordered; therefore, we can use the magnetization $M$ defined below as an order parameter to characterize the ground state,
@@
where $S^{(i)}_{\alpha}$ is the $\alpha$-component of the spin operator defined on the $i$-th inequivalent site in the unit cell, $|m_{i}|$ is the length of the magnetic vector $m_{i}$, and $N$ is the total number of inequivalent sites in a unit cell. The obtained magnetization $M$ as a function of $D$ is plotted in Fig. 6. Similarly to $E_{g}$, the magnetization is also systematically lowered as $D$ increases. When $D=9$, $M$ is lowered to 0.31897, and the extrapolated value in the large-$D$ limit obtained by a power fitting of the data $M(1/D)$ gives an estimation $M=0.3159$, which is only slightly higher than the value $M=0.307$ in quantum Monte Carlo simulation HM-MC , but comparable to those from double-layer and variational update calculations LW-PTNS2019 ; Corboz-vf . What’s more, we find that the local magnetization $m_{i}$ exhibits an explicit antiferromagnetic pattern indicated by staggered orientations. All these signatures are consistent with the Néel antiferromagnetic ground state.
Figure 6: The obtained magnetization $M$ of the square lattice Heisenberg model. A power-law fitting of data (dashed line) gives $M=0.3159$ in the large-$D$ limit. The solid red line marks the value $M=0.307$ from a quantum Monte Carlo estimation HM-MC . The solid black and green lines mark the values of $M$ from a $D=7$ double layer LW-PTNS2019 and a $D=6$ variational update Corboz-vf tensor network calculations.
As described in Sec. II, there are two important hyperparameters in variational tensor network calculations. One is the bond dimension $D$ of the PEPS ansatz as expressed in Eq. (LABEL:Eq:PEPS), the other is the environment dimension $\chi$ used in the CTMRG algorithms to evaluate $\langle\Psi|\Psi\rangle$ for a given $D$. Essentially, the calculation is variational with respect to $D$, but is not variational for $\chi$. Therefore, for a given $D$, the expectation values $E(D)$ and $M(D)$ are reliable only under the condition that they are converged with respect to $\chi$ as $\chi$ increases NTN2017 . Therefore, in this work, in order to ensure the convergence of $E$ and $M$ for each $D$, we always keep the environment dimension $\chi$ sufficiently large. A detailed analysis for $D=9$ is shown in Fig. 7. It shows that by increasing $\chi$ to 160, both the energy and the magnetization are systematically improved, and the convergence error is only about the order of $10^{-6}$, which is already rather satisfactory for our purpose.
@@
III.2 The frustrated Shatry-Sutherland model
@@
where $\boldsymbol{S}_{i}$ is the three-component spin operator defined on $i$-the lattice point. A sketch of the Shastry-Sutherland lattice is illustrated in Fig. 8(a). From the view of the underlying square lattice, one can see that in Eq. (11), $J$ denotes the nearest neighboring antiferromagnetic coupling, and $J^{\prime}$ denotes the next nearest neighboring coupling defined in only one set of squares and arranged as a wheelchair pattern. This model draws attentions in recent years Corboz-SSM ; Corboz-SSMH ; NingXi-SSM2023 ; WYL-PRL2024 because of the stabilization of a quantum disordered plaquette valence bond solid ground state in the intermediate $J/J^{\prime}$ regime of the phase diagram, as well as possible realization of a deconfined quantum critical point DQCP-Senthil and/or quantum spin liquid Corboz-QSLSSM2025 ; WYL-PRL2024 ; LingWang-SSM relevant to some interesting experimental results in real materials SCBO-Science .
@@
Figure 8: The Shastry-Sutherland model we studied in this work, as expressed in Eq. (11). (a) The Shasty-Sutherland lattice. The grey bonds constitute the underlying square lattice, and the coupling strength on these bonds corresponds to $J$. The black bonds correspond to the coupling $J^{\prime}$ that forms a wheelchair pattern. (b) A sketch of the phase diagram consisting of a dimer singlet, a plaquette valence bond solid (VBS), and a Néel antiferromagnetic ground states with varying $J/J^{\prime}$. The nature of the plaquette VBS to the antiferromagnetic transition is under debate, with the possiblity of an intervening quantum spin liquid as marked by the red dashed region. (c) The bond correlation $\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}\rangle$ in the obtained ground state at $J/J^{\prime}=0.7$ in this work, consistent with the empty-plaquette VBS state NingXi-SSM2023 .
@@
An illustration of the phase diagram is also shown in Fig. 8(b). When $J$ dominates, the ground state of this model is a Néel antiferromagnetic state, while when $J^{\prime}$ dominates, the two spins coupled by $J^{\prime}$, as indicated by the black bonds in Fig. 8(a), forming a dimer singlet, and the ground state of the system is a direct product of these dimer singlets. In the intermediate regime, the ground state is a plaquette valence bond solid. However, there is much controversy on the competition between the full-plaquette and empty-plaquette VBS patterns Corboz-SSM ; NingXi-SSM2023 ; SCBO-Science , the nature of the phase transition between the Néel and VBS phases NingXi-SSM2023 ; SCBO-Science , and the possible existence of an intervening quantum spin liquid regime Corboz-QSLSSM2025 ; WYL-PRL2024 ; LingWang-SSM . In this work, using the proposed single-layer variational framework and setting $J^{\prime}=1$, we focus on two parameter settings: the possible VBS phase with $J=0.7$, and the antiferromagnetic phase with $J=0.8$.
@@
In Fig. 9, we focus on $J=0.7$, and plot the obtained ground-state energy $E_{g}$ as a function of $D$. The variational feature remains pronounced, i.e., the energy systematically improves as $D$ becomes larger. When $D=9$, we have $E_{g}=-0.3880225$. If we fit the data $E_{g}(1/D)$ and extrapolate to the large-$D$ limit, we obtain an estimation $E_{g}=-0.3885$ black dashed in Fig. 9(a)), which is consistent with the previous PEPS from finite systems WYL-PRL2024 . To make the result convincing, we analyze the energy for $D=9$ with $\chi$ ranging from 90 to 140 in Fig. 9(b). It is clear that the energy has almost fully converged with a precision of about $10^{-6}$.
@@
Figure 9: The obtained ground-state energy $E_{g}$ of the Shastry-Sutherland model at $J=0.7$. (a) $E_{g}$ as a function of bond dimension $D$, with $D$ ranging from 2 to 9. A power-law fitting of data $E_{g}(1/D)$ is indicated by a black dashed line. For comparison, an extrapolation from finite PEPS calculations WYL-PRL2024 is also plotted as a red solid line. (b) The convergence analysis of $E_{g}(\chi)$ for $D=9$.
@@
Further analysis shows that the obtained ground state is an empty-plaquette VBS state, which is illustrated in Fig. 8(c). To show this, firstly we introduce an order parameter $M_{p}$ defined in the following
$M_{p}=\frac{1}{4}\Big|\sum_{\langle i,j\rangle\in\square}\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle-\sum_{\langle i,j\rangle\in\boxslash}\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle\Big|$ (12)
[…]
@@
For $J=0.8$, we plot the ground-state energy in Fig. 11. The behavior is similar to that for $J=0.7$, and the energy systematically lowers and gradually converges as $D$ becomes larger. When $D=9$, we have $E_{g}=-0.448708$. This is already very close to the previous tensor network studies, such as an infinite projected entangled simplex state (iPESS) study NingXi-SSM2023 , an infinite PEPS study employing U(1) spin symmetry Corboz-QSLSSM2025 , and an extrapolation from finite PEPS calculation WYL-PRL2024 . These estimations are also included in Fig. 11 for comparison. Extrapolation to the large-$D$ limit via a simple power-law fitting yields an estimate $E_{g}=-0.4490$, which is more reliable. With $\chi$ up to 160, Fig. 11 shows that the convergence precision of the ground-state energy for $D=9$ has reached the order $10^{-6}$, indicating the obtained result is convincing.
Figure 11: The obtained ground-state energy $E_{g}$ of the Shastry-Sutherland model at $J=0.8$. (a) $E_{g}$ as a function of bond dimension $D$, with $D$ ranging from 2 to 9. A power-law fitting of data $E_{g}(1/D)$ is indicated by a black dashed line. For comparison, an iPESS study NingXi-SSM2023 , an infinite PEPS study employing U(1) symmetry Corboz-QSLSSM2025 , and an extrapolation from finite PEPS calculations WYL-PRL2024 , are also included as solid lines. (b) The convergence analysis of $E_{g}(\chi)$ for $D=9$.
In Fig. 12, we plot the magnetization $M$ of the obtained ground state as a function of $D$. It shows that though $M$ becomes weaker and weaker as $D$ increases, it remains finite and nonzero. For example, when $D=9$, we still obtain a significant magnetization of about 0.1303, and an extrapolation in the large-$D$ limit still gives a nonzero estimation of about $M=0.12$. This is neither an artificial finite-$\chi$ effect, since by pushing $\chi$ to 160, the convergence precision of $M$ for $D=9$ has also reached the order $10^{-6}$, as clearly shown in Fig. 12(b). This is unambiguously consistent with the fact that $J=0.8$ falls in the antiferromagnetic phase.
@@ IV Discussions
In this work, we propose a single-layer framework that integrates the nested tensor network method into the variational tensor network calculations leveraging automatic differentiation. This framework can reduce the memory and computational costs by two and three orders of $D$ compared to traditional variational calculations, thereby enabling $D$ to be larger and achieving better accuracy. The validity of the framework is demonstrated through two prototypical models: the square lattice Heisenberg model and the Shastry-Sutherland model. Even without GPU hardware acceleration or symmetry considerations, we can extend $D$ to 9, yielding accurate ground-state energies and consistent order parameters with previous studies. Especially, for the Shastry-Sutherland model, we confirm an intermediate empty-plaquette VBS phase in the ground state diagram. The convergence with respect to $D$ and $\chi$ is systematically analyzed, and nearly complete convergence is observed even for the largest $D$. We believe this framework could enable large-scale variational tensor network calculations in more complex quantum systems, such as superconductivity TNStJ ; TNSHubbard , excited states, and dynamical studies ADPEPS ; LY-PRB2025 in the future.
@@
In our calculations, the single-layer tensor network with nested structure is contracted by the CTMRG algorithm. It is known that the idea of nested tensor network can be performed not only in CTMRG YJKao-CTMRG ; DNSheng-CTMRG , but also in other tensor contraction algorithms, such as time-evolving block decimation NTN2017 ; Vidal-TEBD ; Orus-TEBD , and coarse-graining renormalization group. Therefore, correspondingly, the single-layer framework in this work can also be realized in other algorithms than CTMRG. The extra acceptable sacrifice is probably the same, i.e., the size of the unit cell of the single-layer nested tensor network may be doubled, as shown in Fig. 2(c) in this work.
@@
The bond dimension $D$ can be further increased by incorporating additional advanced techniques into this framework. For example, in this work we did not consider the spin symmetry, but one can leverage the U(1) symmetry of the two models to improve efficiency. For the square lattice Heisenberg model in this work, one can utilize a local spin rotation $i\sigma_{y}$ on one set of lattice locations Hasik-Sci2021 ; LY-PRB2025 to reduce the $4\times 4$ unit cell to a $2\times 2$ unit cell in the generated nested tensor network, which can simplify the calculations shown in Fig. 2(c). Even for the Shastry-Sutherland model, this strategy can accelerate the convergence in variational calculations. Moreover, one can use a checkpoint LW-PTNS2019 or a fixed-point technique to reduce the length of the chain, i.e., Eq. (5) used in the back propagation, by calculating the gradient from a given or an approximate fixed-point environment LW-fixedMPS . The latter strategy is similar to the finite-system approximation SRG2010 or to choosing a better initial value for the CTMRG environment and can effectively reduce memory and computational costs in automatic differentiation calculations. Finally, at a purely computational level, one can use partial singular value decomposition, randomized decomposition RSVD in Eq. (9), or employ a more economical QR decomposition Corboz-QR2025 instead, to reduce computational cost further.
@@
It is worth noting that several recent studies also involve single-layer tensor network states in the literature. For example, in Ref. Schmoll-PRB2025 , Jan et al. proposed a similar framework called the split-CTMRG algorithm, which is also based on the nested tensor network structure revealed in Ref. NTN2017 . However, the contraction scheme is designed particularly and differs from what we proposed in Sec. II.3. In Ref. ZGM-PRB2025 , Xu et al. extended the idea of nested structure to a triple-layer tensor network and applied it to the three-dimensional Ising model successfully, while earlier in Ref. YLP-PRB2023 , Yang et al. proposed a new nesting structure which is more natural and symmetric than the one proposed in Ref. NTN2017 and was applied successfully to the same model. Probably, these single-layer proposals can be further improved by combining their distinct advantages, which we would like to leave as a future pursuit.

### v2

@@ III.1 The Heisenberg model on a square lattice
Figure 4: The PEPS ansatz used in this work to study the square-lattice Heisenberg model. Dashed lines separate the $2\times 2$ unit cells.
The estimated ground-state energy is plotted in Fig. 5 as a function of $D$. It shows that the obtained energy systematically decreases as $D$ increases, verifying the variational nature of the framework. When $D=9$, we obtain an estimate of $E_{g}=-0.6693896$, which is already very close to the value $E_{g}=-0.6694421$ in quantum Monte Carlo simulation [35]. If we make a further power-law fitting of $E_{g}(1/D)$, then in the large-$D$ limit we obtain an extrapolated value of $E_{g}=-0.66941$, which deviates from the quantum Monte Carlo estimate only by the order of $10^{-5}$.
Figure 5: The obtained ground-state energy $E_{g}$ of the square lattice Heisenberg model. (a) $E_{g}$ as a function of bond dimension $D$, with $D$ ranging from 2 to 9. (b) A power-law fitting of data $E_{g}(1/D)$. The extrapolation to the large-$D$ limit gives $E_{g}=-0.66941$. A quantum Monte Carlo estimate from Ref. [35], $E_{g}=-0.6694421$, is also shown as red solid.
As mentioned before, the ground state of this model is antiferromagnetically long-range ordered; therefore, we can use the magnetization $M$ defined below as an order parameter to characterize the ground state:
@@
where $S^{(i)}_{\alpha}$ is the $\alpha$ component of the spin operator defined on the $i$ th inequivalent site in the unit cell, $|m_{i}|$ is the length of the magnetic vector $m_{i}$, and $N$ is the total number of inequivalent sites in a unit cell. The obtained magnetization $M$ as a function of $D$ is plotted in Fig. 6. Similarly to $E_{g}$, the magnetization is also systematically lowered as $D$ increases. When $D=9$, $M$ is lowered to 0.31897, and the extrapolated value in the large-$D$ limit obtained by a power fitting of the data $M(1/D)$ gives an estimate $M=0.3159$, which is only slightly higher than the value $M=0.307$ in quantum Monte Carlo simulation [35], but comparable to those from double-layer and variational update calculations [27, 45]. What is more, we find that the local magnetization $m_{i}$ exhibits an explicit antiferromagnetic pattern indicated by staggered orientations. All these signatures are consistent with the Néel antiferromagnetic ground state.
Figure 6: The obtained magnetization $M$ of the square lattice Heisenberg model. A power-law fitting of data (dashed line) gives $M=0.3159$ in the large-$D$ limit. The solid red line marks the value $M=0.307$ from a quantum Monte Carlo estimate [35]. The solid black and green lines mark the values of $M$ from a $D=7$ double layer [27] and a $D=6$ variational update [45] tensor network calculations.
As described in Sec. II, there are two important hyperparameters in variational tensor network calculations. One is the bond dimension $D$ of the PEPS ansatz as expressed in Eq. (LABEL:Eq:PEPS), and the other is the environment dimension $\chi$ used in the CTMRG algorithms to evaluate $\langle\Psi|\Psi\rangle$ for a given $D$. Essentially, the calculation is variational with respect to $D$, but is not variational for $\chi$. Therefore, for a given $D$, the expectation values $E(D)$ and $M(D)$ are reliable only under the condition that they are converged with respect to $\chi$ as $\chi$ increases [34]. Therefore, in this work, in order to ensure the convergence of $E$ and $M$ for each $D$, we always keep the environment dimension $\chi$ sufficiently large. A detailed analysis for $D=9$ is shown in Fig. 7. It shows that by increasing $\chi$ to 160, both the energy and the magnetization are systematically improved, and the convergence error is only of the order of $10^{-6}$, which is already rather satisfactory for our purpose.
@@
III.2 The frustrated Shastry-Sutherland model
@@
where $\boldsymbol{S}_{i}$ is the three-component spin operator defined on $i$ th lattice point. A sketch of the Shastry-Sutherland lattice is illustrated in Fig. 8(a). From the view of the underlying square lattice, one can see that in Eq. (11), $J$ denotes the nearest-neighboring antiferromagnetic coupling, and $J^{\prime}$ denotes the next-nearest-neighbor coupling defined in only one set of squares and arranged as a wheelchair pattern. This model draws attention in recent years [36, 46, 10, 47] because of the stabilization of a quantum disordered plaquette valence bond solid ground state in the intermediate $J/J^{\prime}$ regime of the phase diagram, as well as possible realization of a deconfined quantum critical point [48] and/or quantum spin liquid [50, 47, 49] relevant to some interesting experimental results in real materials [51].
@@
Figure 8: The Shastry-Sutherland model we studied in this work, as expressed in Eq. (11). (a) The Shasty-Sutherland lattice. The grey bonds constitute the underlying square lattice, and the coupling strength on these bonds corresponds to $J$. The black bonds correspond to the coupling $J^{\prime}$ that forms a wheelchair pattern. (b) A sketch of the phase diagram consisting of a dimer singlet, a plaquette valence bond solid (VBS), and a Néel antiferromagnetic ground state with varying $J/J^{\prime}$. The nature of the plaquette VBS to the antiferromagnetic transition is under debate, with the possiblity of an intervening quantum spin liquid as marked by the red dashed region. (c) The bond correlation $\langle\boldsymbol{S}_{i}\cdot\boldsymbol{S}_{j}\rangle$ in the obtained ground state at $J/J^{\prime}=0.7$ in this work, consistent with the empty-plaquette VBS state [10].
@@
An illustration of the phase diagram is also shown in Fig. 8(b). When $J$ dominates, the ground state of this model is a Néel antiferromagnetic state, while when $J^{\prime}$ dominates, the two spins coupled by $J^{\prime}$, as indicated by the black bonds in Fig. 8(a), form a dimer singlet, and the ground state of the system is a direct product of these dimer singlets. In the intermediate regime, the ground state is a plaquette valence bond solid. However, there is much controversy over the competition between the full-plaquette and empty-plaquette VBS patterns [36, 10, 51], the nature of the phase transition between the Néel and VBS phases [10, 51], and the possible existence of an intervening quantum spin liquid regime [50, 47, 49]. In this work, using the proposed single-layer variational framework and setting $J^{\prime}=1$, we focus on two parameter settings: the possible VBS phase with $J=0.7$, and the antiferromagnetic phase with $J=0.8$.
@@
In Fig. 9, we focus on $J=0.7$, and plot the obtained ground-state energy $E_{g}$ as a function of $D$. The variational feature remains pronounced, i.e., the energy systematically improves as $D$ becomes larger. When $D=9$, we have $E_{g}=-0.3880225$. If we fit the data $E_{g}(1/D)$ and extrapolate to the large-$D$ limit, we obtain an estimate $E_{g}=-0.3885$ [black dashed line in Fig. 9(a)], which is consistent with the previous PEPS from finite systems [47]. To make the result convincing, we analyze the energy for $D=9$ with $\chi$ ranging from 90 to 140 in Fig. 9(b). It is clear that the energy has almost fully converged with a precision of about $10^{-6}$.
@@
Figure 9: The obtained ground-state energy $E_{g}$ of the Shastry-Sutherland model at $J=0.7$. (a) $E_{g}$ as a function of bond dimension $D$, with $D$ ranging from 2 to 9. A power-law fitting of data $E_{g}(1/D)$ is indicated by a black dashed line. For comparison, an extrapolation from finite PEPS calculations [47] is also plotted as a red solid line. (b) The convergence analysis of $E_{g}(\chi)$ for $D=9$.
@@
Further analysis shows that the obtained ground state is an empty-plaquette VBS state, which is illustrated in Fig. 8(c). To show this, first we introduce an order parameter $M_{p}$ defined as
$M_{p}=\frac{1}{4}\Big|\sum_{\langle i,j\rangle\in\square_{A}}\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle-\sum_{\langle i,j\rangle\in\square_{B}}\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle\Big|$ (12)
where the two $\sum$ sum over all the nearest-neighbor bond energies in inequivalent empty squares $\square_{A}$ and $\square_{B}$ [olive and red plaquettes illustrated in Fig. 8(c)]. In Fig. 10(a), we plot the obtained $M_{p}$ as a function of $D$. It shows clearly that the order parameter converges gradually as $D$ increases, and when $D=9$ we obtain $M_{p}=$ 0.1945. By extrapolating to the large-$D$ limit, we obtain a final estimate $M_{p}=$ 0.1930, which is a strong indication that the ground state breaks the $Z_{2}$ symmetry between two inequivalent squares. The convergence with respect to $\chi$ plotted in Fig. 10(b) indicates that the symmetry breaking is robust and is not a finite-$\chi$ effect. To show the detailed pattern and distinguish the two possible VBS states [10, 36], we plot the bond correlation $\langle\mathbf{S}_{i}\cdot\mathbf{S}_{j}\rangle$ for each bond in Fig. 8(c). It shows that the correlation in squares without $J^{\prime}$ couplings is much stronger than that in squares with $J^{\prime}$ couplings, consistent with the empty-plaquette VBS state. Furthermore, we find the obtained ground state has a tiny magnetization of $\sim 10^{-4}$, suggesting it probably does not break the SU(2) spin symmetry. Combining these features, we can convincingly conclude that the ground state is an empty-plaquette VBS state at $J=0.7$, consistent with previous studies [10].
@@
For $J=0.8$, we plot the ground-state energy in Fig. 11. The behavior is similar to that for $J=0.7$, and the energy systematically decreases and gradually converges as $D$ becomes larger. When $D=9$, we have $E_{g}=-0.448708$. This is already very close to the previous tensor network studies, such as an infinite projected entangled simplex state (iPESS) study [10], an infinite PEPS study employing U(1) spin symmetry [50], a neural network study [52] and an extrapolation from finite PEPS calculation [47]. These estimate are also included in Fig. 11 for comparison. Extrapolation to the large-$D$ limit via a simple power-law fitting yields an estimate of $E_{g}=-0.4490$, which is more reliable. With $\chi$ up to 160, Fig. 11 shows that the convergence precision of the ground-state energy for $D=9$ has reached the order $10^{-6}$, indicating that the obtained result is convincing.
Figure 11: The obtained ground-state energy $E_{g}$ of the Shastry-Sutherland model at $J=0.8$. (a) $E_{g}$ as a function of bond dimension $D$, with $D$ ranging from 2 to 9. A power-law fitting of data $E_{g}(1/D)$ is indicated by a black dashed line. For comparison, an iPESS study [10], an infinite PEPS study employing U(1) symmetry [50], a neural network study [52] and an extrapolation from finite PEPS calculations [47], are also included as solid lines. (b) The convergence analysis of $E_{g}(\chi)$ for $D=9$.
In Fig. 12, we plot the magnetization $M$ of the obtained ground state as a function of $D$. It shows that though $M$ becomes weaker and weaker as $D$ increases, it remains finite and nonzero. For example, when $D=9$, we still obtain a significant magnetization of about 0.1303, and an extrapolation in the large-$D$ limit still gives a nonzero estimate of about $M=0.12$. This is neither an artificial finite-$\chi$ effect, since by pushing $\chi$ to 160, the convergence precision of $M$ for $D=9$ has also reached the order $10^{-6}$, as clearly shown in Fig. 12(b). This is unambiguously consistent with the fact that $J=0.8$ falls in the antiferromagnetic phase.
@@ IV Discussions
In this work, we propose a single-layer framework that integrates the nested tensor network method into the variational tensor network calculations leveraging automatic differentiation. This framework can reduce the memory and computational costs by two and three orders of $D$ compared to traditional variational calculations, thereby enabling $D$ to be larger and achieving better accuracy. The validity of the framework is demonstrated through two prototypical models: the square-lattice Heisenberg model and the Shastry-Sutherland model. Even without GPU hardware acceleration or symmetry considerations, we can extend $D$ to 9, yielding accurate ground-state energies and consistent order parameters with previous studies. In particular, for the Shastry-Sutherland model, we confirm an intermediate empty-plaquette VBS phase in the ground-state phase diagram. The convergence with respect to $D$ and $\chi$ is systematically analyzed, and nearly complete convergence is observed even for the largest $D$. We believe this framework could enable large-scale variational tensor network calculations in more complex quantum systems, such as superconductivity [12, 13], excited states, and dynamical studies [53, 11] in the future.
@@
In our calculations, the single-layer tensor network with nested structure is contracted by the CTMRG algorithm. It is known that the idea of nested tensor network can be performed not only in CTMRG [40, 41], but also in other tensor contraction algorithms, such as time-evolving block decimation [34, 42, 43], and coarse-graining renormalization group. Therefore, correspondingly, the single-layer framework in this work can also be realized in other algorithms than CTMRG. The extra acceptable sacrifice is probably the same, i.e., the size of the unit cell of the single-layer nested tensor network may be doubled, as shown in Fig. 2(c) in this work.
@@
The bond dimension $D$ can be further increased by incorporating additional advanced techniques into this framework. For example, in this work we did not consider the spin symmetry, but one can leverage the U(1) symmetry of the two models to improve efficiency. For the square lattice Heisenberg model in this work, one can utilize a local spin rotation $i\sigma_{y}$ on one set of lattice locations [54, 11] to reduce the $4\times 4$ unit cell to a $2\times 2$ unit cell in the generated nested tensor network, which can simplify the calculations shown in Fig. 2(c). Even for the Shastry-Sutherland model, this strategy can accelerate the convergence in variational calculations. Moreover, one can use a checkpoint [27] or a fixed-point technique to reduce the length of the chain, i.e., Eq. (5) used in the back propagation, by calculating the gradient from a given or an approximate fixed-point environment [55]. The latter strategy is similar to the finite-system approximation [56] or to choosing a better initial value for the CTMRG environment and can effectively reduce memory and computational costs in automatic differentiation calculations. Finally, at a purely computational level, one can use partial singular value decomposition, randomized decomposition [57] in Eq. (9), or employ a more economical QR decomposition [58] instead, to reduce computational cost further.
As reviewed in Sec. II.2, the nested tensor network method treats the bra $\langle\Psi|$ and ket $|\Psi\rangle$ independently when contracting $\langle\Psi|\Psi\rangle$, thus the Hermiticity of the environment (for the reduced tensor) [59] can be easily broken. Fortunately, the Hermiticity can be gradually restored by increasing the environment dimension $\chi$, and could be further improved by considering the hidden $Z_{2}$ symmetry [59], more delicately designed iterations [60], and smarter variation with better convergence [61], etc. The restoration of the Hermiticity is expected to improve the accuracy of the results.
@@
It is worth noting that several recent studies also involve single-layer tensor network states in the literature. For example, in Ref. [60], Jan et al. proposed a similar framework called the split-CTMRG algorithm, which is also based on the nested tensor network structure revealed in Ref. [34]. However, the contraction scheme is particularly designed and differs from what we proposed in Sec. II.3. In Ref. [62], Xu et al. extended the idea of the nested structure to a triple-layer tensor network and applied it to the three-dimensional Ising model successfully, while earlier in Ref. [63], Yang et al. proposed a new nesting structure that is more natural and symmetric than the one proposed in Ref. [34] and was applied successfully to the same model. Probably, these single-layer proposals can be further improved by combining their distinct advantages, which we would like to leave as a future direction.